OpenStudy (anonymous):

FAN AND MEDAL!!! A line passes through the point four comma negative one and has a slope of three-halves. If you use this slope to generate another point on the line, what are the coordinates of that point? eight comma negative three six comma two twelve comma negative two seven comma one

OpenStudy (mrnood):

 The formula for a straight line is y=mx +c where m is the slope You know 1 point (4,-1)and the slope 3/2 so -1= 3x/2 +c Use this to calculate c you then have a formula for the line y=3x/2 + c Try the other pairs of numbers to see which pair give the correct solution to this formula...

OpenStudy (anonymous):

so what is it

OpenStudy (mrnood):

I am not going to do the sum for you the method is given above If you correctly calculate c then I will give more help if you need it

OpenStudy (anonymous):

I don't get it

OpenStudy (mrnood):

Are you studying the equation of a straight line?

OpenStudy (mrnood):

If you want help please answer

OpenStudy (anonymous):

i need help i don't understand

OpenStudy (mrnood):

do you know the formula I gave above y=mx+c for a straight line?

OpenStudy (anonymous):

your no help

OpenStudy (mrnood):

I CAN be of help - but I am NOT going to give you the answer if that is what oyu mean. I am her giving my time to help you understand. NOT to do sums which I can do in my head

OpenStudy (mrnood):

AND by the way - you mean "You're no help" not "Your no help"